show that only one out of n, n+2 or n+4 is divisible by 3 where n is positive integer.
Answer by Edwin McCravy(20086)

show that only one out of n, n+2 or n+4 is divisible by 3 where n is positive integer.

By induction:

When n = 1, exactly one of 1, 1+2, and 1+4 is divisible by 3,
namely 1+2, since 3 is divisible by 3, and the other two 1 and
5 are not.

Suppose for n < k, only one out of n, n+2, n+4 is divisible by 3

For n = k, we consider k, k+2, k+4.  

By the induction hypothesis, only one of k-1, k+1 and k+3 is 
divisible by 3.  We look at the three possible cases.

Case 1: k-1 is the one which is divisible by 3. Then k-1 = 3m, for
some positive integer m.  

Then add 1 to both sides of k-1 = 3m:

k-1+1=3m+1
k = 3m+1

then k is NOT divisible by 3

Now add 3 to both sides of k-1 = 3m:

k-1+3=3m+3
k+2 = 3m+3
k+2 = 3(m+1)

then k+2 IS divisible by 3

Now add 5 to both sides of k-1 = 3m:

k-1+5=3m+5
k+4 = 3m+5
k+4 = 3m+3 + 2 = 3(m+1)+2

then k+4 is NOT divisible by 3.

So, we have proved case 1 for n = k

------------------------

Case 2: k+1 is the one which is divisible by 3. Then k+1 = 3m, for
some positive integer m.  

Then add -1 to both sides of k+1 = 3m:

k+1-1=3m-1
k = 3m-1

then k is NOT divisible by 3

Now add 1 to both sides of k+1 = 3m:

k+1+1=3m+1
k+2 = 3m+1

then k+2 is NOT divisible by 3

Now add 3 to both sides of k+1 = 3m:

k+1+3=3m+3
k+4 = 3m+3
k+4 = 3(m+1)

so k+4 IS divisible by 3.

So, we have proved case 2.

------------------------ 

Case 3: k+3 is the one which is divisible by 3. Then k+3 = 3m, for
some positive integer m.  

Then add -3 to both sides of k+3 = 3m:

k+3-3=3m-3
k = 3(m-1)

then k IS divisible by 3

Now add -1 to both sides of k+3 = 3m:

k+3-1=3m-1
k+2 = 3m-1

then k+2 is NOT divisible by 3

Now add 1 to both sides of k+3 = 3m:

k+3+1=3m+1
k+4 = 3m+1

then k+4 is NOT divisible by 3.

So, we have proved case 3.

The proof is now complete.

Edwin